# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 5.2. Truncated categories
We have $\mathrm{Ext}^{i}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0$ only if $x\leq^{\frac{\infty}{2}}x^{\prime}$.
Consider the induction functor
$\mathrm{ind}_{c}:=({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}-:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}.$
Note that $\mathrm{ind}_{c}(M)=U{\mathfrak{n}}\otimes M$ as vector spaces, for any ${\mathrm{A}}^{b}_{S}$-module $M$.
@@
For any objects $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, the induction $\mathrm{ind}_{c}(P)$ also admits Verma flags.
@@
It is enough to prove the assertion when $P$ is a Verma object. By Corollary 4.11, $P$ is a module of ${\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}}$. Consider the homomorphism of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules
(5.2) $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}P=U{\mathfrak{n}}\otimes P\rightarrow P\otimes U{\mathfrak{n}},$
$u\otimes m\mapsto u^{(1)}.m\otimes u^{(2)},\quad\forall u\in U{\mathfrak{n}},\ \forall m\in P,$
where $P\otimes U{\mathfrak{n}}$ is viewed as an $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-module by ${\mathrm{A}}^{b}_{S}$ acting on $P$ and $U{\mathfrak{n}}$ acting on $P\otimes U{\mathfrak{n}}$ diagonally. The homomorphism (5.2) is an isomorphism, whose inverse is given by $m\otimes u\mapsto u^{(2)}\otimes(\mathrm{S}u^{(1)}).m$ ($\mathrm{S}$ is the antipode for $U{\mathfrak{n}}$). We obtain an isomorphism $\mathrm{ind}_{c}(P)\simeq P\otimes U{\mathfrak{n}}$ of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules. It is clear that $P\otimes U{\mathfrak{n}}$ admits a Verma flag with Verma factors $P\otimes{\mathbb{C}}_{\eta}$ for $\eta\geq 0$. ∎
Note the subset $\{t(\lambda)w\}_{w\in W,\lambda\leq\nu}$ in an ideal in $(W_{\mathrm{ex}},\leq^{\frac{\infty}{2}})$. By the lemma above and Proposition 5.5, for any object $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, one can define the quotients
$\mathrm{ind}_{c}(P)^{\leq\nu},\quad\nu\in\Lambda$
of $\mathrm{ind}_{c}(P)$ modulo the submodule with Verma factors in $\{{\mathcal{M}}_{t(\lambda)w}\}_{w\in W,\ \lambda\nleq\nu}$. Then the $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules $\mathrm{ind}_{c}(P)^{\leq\nu}$ are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
@@
The family $\{\mathrm{ind}_{c}(P)^{\leq\nu}\}$ with $P$ running over a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
@@
Let $x=t(\lambda)w$ where $\lambda\in\Lambda$ and $w\in W$, then the truncation $M(x\bullet_{l}0)^{\leq l\nu}_{S}\neq 0$ only if $\lambda\leq\nu$. Hence the functor $\mathrm{Hom}({\mathcal{M}}_{x},-)$ is nonzero on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$ only if $\lambda\leq\nu$. It shows that any homomorphism from $\mathrm{ind}_{c}(P)$ to an object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ factors through the quotient $\mathrm{ind}_{c}(P)^{\leq\nu}$. It follows natural isomorphisms
(5.3) $\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}(\mathrm{ind}_{c}(P)^{\leq\nu},-)$ $=\mathrm{Hom}_{({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}}(\mathrm{ind}_{c}(P),-)$
$=\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}}(P,-)$
of functors on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. Hence $\mathrm{ind}_{c}(P)^{\leq\nu}$ is projective in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. On the other hand, for any module $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, there exists a surjection of $\Lambda$-graded ${\mathrm{A}}^{b}_{S}$-modules $P\twoheadrightarrow M$ for some projective $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Then the induced homomorphism $\mathrm{ind}_{c}(P)\twoheadrightarrow\mathrm{ind}_{c}(P)^{\leq\nu}\rightarrow M$ is a surjection. So we obtain a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. ∎
@@
Any $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a maximal quotient $M^{\leq\nu}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. It defines a functor
$(-)^{\leq\nu}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}\rightarrow{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$
left adjoint to the inclusion functor. We have a natural isomorphism
${\mathbb{V}}_{c}\circ(-)^{\leq\nu}=(-)^{\leq\nu}\circ{\mathbb{V}}_{c}.$
@@
We may assume $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu^{\prime}}$ for some $\nu^{\prime}>>\nu$. By Lemma 5.7, we can choose a presentation $\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}}\rightarrow M\rightarrow 0$, for some projective object $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. It induces a homomorphism $\mathrm{ind}_{c}(P_{1})^{\leq\nu}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu}$, and we set $M^{\leq\nu}$ to be its cokernel. Then $M^{\leq\nu}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. For any $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, we have a commutative diagram with exact rows
00{0}${\mathrm{Hom}(M,M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}},M^{\prime})}$00{0}${\mathrm{Hom}(M^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu},M^{\prime}),}$
where the two vertical equalities follow from the first equality in (5.3). Hence the left vertical map is an isomorphism. It shows that $M^{\leq\nu}$ is the maximal quotient of $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, and therefore $(-)^{\leq\nu}$ defines a functor left adjoint to the inclusion.
For the last assertion, by the right exactness of ${\mathbb{V}}_{c}\circ(-)^{\leq\nu}$ and $(-)^{\leq\nu}\circ{\mathbb{V}}_{c}$, it is enough to construct natural isomorphism ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$ for any projective module $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. By Lemma 5.4, the module ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))$ admits a filtration with composition factors given by ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ for some $\lambda\in\Lambda$ and $w\in W$. By weight consideration, we have
$\mathrm{Ext}^{1}({\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda},{\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w^{\prime}}S\otimes{\mathbb{C}}_{\mu})\neq 0\quad\text{only if}\quad\lambda<\mu,\quad\forall w,w^{\prime}\in W.$
Note that ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})$ is the quotient by the submodule extended by the factors ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ with $\lambda\nleq\nu$, which is exactly the truncation ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$. ∎
We have a parallel result of Lemma 5.8.
Lemma 5.9.
There is a natural isomorphism
${\mathbb{V}}_{r}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ{\mathbb{V}}_{r}$
Proof.
By [36, Lem 4.3(3)] we have a natural isomorphism $T^{-\rho}_{0}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ T^{-\rho}_{0}$, which upgrades to the desired natural isomorphism. ∎
@@
Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$. Since the restriction of $\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ on $\{e\}\times{\mathfrak{b}}\hookrightarrow\tilde{{\mathfrak{g}}}$ is ${\mathrm{A}}^{b}_{S}$, we can identify
(5.4) ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$
by induction from $B$ to $G$. The ${\mathrm{A}}^{b}_{S}$-module ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ corresponds to ${\mathcal{E}}^{\vee}_{\hat{0}}:={\mathcal{E}}^{\vee}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}$ under (5.4). Let $\mathrm{St}=\tilde{{\mathfrak{g}}}\times_{{\mathfrak{g}}}\tilde{{\mathfrak{g}}}$ be the Steinberg variety, and let $\widehat{\mathrm{St}}$ be its completion at $0\in{\mathfrak{t}}/W$. Let $\pi_{2}:\widehat{\mathrm{St}}\rightarrow\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ be the projection to the second factor. There is an equivalence
$\pi_{2}^{*}{\mathcal{E}}_{{\hat{0}}}\otimes^{L}_{\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}}-:\ D^{b}\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}}),$
sending ${\mathcal{E}}^{\vee}_{\hat{0}}$ to the structure sheaf ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ of the diagonal $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ in $\widehat{\mathrm{St}}$.

### v4

@@ 5.2.3. Truncation by $\leq^{\frac{\infty}{2}}$
We have
(5.4) $\mathrm{Ext}^{i}_{{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0\quad\text{only if}\quad x\leq^{\frac{\infty}{2}}x^{\prime}.$
For a module $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag, by (5.4) we can form the quotient
$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$
of $Q$ by the submodule that is composed by the Verma factors ${\mathcal{M}}_{x}$ with $x\nleq^{\frac{\infty}{2}}t(\nu)$. We define
${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$
as the the full subcategory of modules in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ that admit a surjection $Q\twoheadrightarrow M$ from a module $Q$ admitting a Verma flag with factors ${\mathcal{M}}_{x}$ with $x\leq^{\frac{\infty}{2}}t(\nu)$. Since $W_{\mathrm{ex}}$ is covered by the poset ideals of the form $\{x\in W_{\mathrm{ex}}|x\leq^{\frac{\infty}{2}}t(\nu)\}$, any object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ is a finite direct sum of objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ for some $\nu\in\Lambda$.
@@
There is a truncation functor
$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}:\ ({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$
that is left adjoint to the natural inclusion.
@@
Since ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$, any morphism from $M\in({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$ to an object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ factors through $\tau^{B,\leq l\nu}M$. Hence it is enough to define the functor
(5.5) $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}.$
Let $Q^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag with factors in $\{M(x\bullet_{l}0)_{S}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$, and let $Q^{\prime}\twoheadrightarrow M^{\prime}$ be a surjection. Let $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$ be a projective object. Then any morphism from $Q$ to $M^{\prime}$ can be lifted to $Q^{\prime}$. By (5.4) any morphism from $Q$ to $Q^{\prime}$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$. Hence any morphism from $Q$ to $M^{\prime}$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$, which then is the maximal quotient of $Q$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
In general, let $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$ and choose a projective resolution $Q_{2}\rightarrow Q_{1}\rightarrow M\rightarrow 0$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$. Then we set
$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M):=\mathrm{coker}\big(\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q_{2})\rightarrow\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q_{1})\big).$
Then $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$ is contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. For any $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$, we have a commutative diagram with exact rows
[…]
Hence the left vertical map is an isomorphism, which shows that $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$ is the maximal quotient of $M$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. It gives the desired functor. ∎
@@
  1. (1)
    The family $\{\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P))\}$ with $P$ running over a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. Moreover, any projective object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admits a Verma flag (with factors in $\{{\mathcal{M}}_{x}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$).
  2. (2)
    The category ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is a Serre subcategory of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
@@
(1) The proof of the first assertion is similar to the one of Lemma 5.15. For the second assertion, for any projective object $P\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, by construction $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P))=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}\circ\tau^{B,\leq l\nu}(\mathrm{ind}_{c}(P))$ admits a Verma flag (with factors in $\{{\mathcal{M}}_{x}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$). Hence any projective object $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is a direct summand of an object admitting a Verma flag. Under the equivalence (4.2), the image of $Q$ is a direct summand of a module admitting a Verma flag in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, which thus admits a Verma flag by a standard argument (see e.g. [20, Lem 2.5]), and so does $Q$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Now the assertion follows from Lemma 5.14.
(2) By definition ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is closed under taking quotient objects. Let $M$ be a subobject of $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. Then the inclusion $M\hookrightarrow M^{\prime}$ factors through the quotient $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$, hence $M=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$. So ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is also closed under taking subobjects.
Now we show that ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is closed under extension. Let $0\rightarrow M_{1}\rightarrow M\rightarrow M_{2}\rightarrow 0$ be a short exact sequence in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ with $M_{1},M_{2}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. Then $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$, and we can choose a surjection $Q\twoheadrightarrow M$ from a projective object $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$. We have short exact sequence $0\rightarrow Q^{\prime}\rightarrow Q\rightarrow\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)\rightarrow 0$, where $Q^{\prime}$ is the subobject of $Q$ composed by Verma factors ${\mathcal{M}}_{x}$ with $x\nleq^{\frac{\infty}{2}}t(\nu)$}. Then $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q^{\prime})=0$, so $\mathrm{Hom}(Q^{\prime},M_{i})=0$ ($i=1,2$). It follows that $\mathrm{Hom}(Q^{\prime},M)=0$. Hence the surjection $Q\twoheadrightarrow M$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)\twoheadrightarrow M$, which implies that $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. ∎
Recall the truncation functor by $\nu\in\Lambda$ (for any commutative Noetherian $S$-algebra $R$)
$\tau^{\leq\nu}_{c}:\ ({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}.$
@@
The functor ${\mathbb{V}}_{c}$ restricts to
${\mathbb{V}}_{c}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}.$
We have a natural isomorphism of functors from $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$ to ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}$,
(5.6) ${\mathbb{V}}_{c}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}=\tau_{c}^{\leq\nu}\circ{\mathbb{V}}_{c}.$
@@
Recall the functor $\pi_{*}:{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{T}$ and its right adjunction $\pi^{!}$. They induces functors (still denoted by $\pi_{*}$ and $\pi^{!}$) for $U{\mathfrak{n}}$-equivariant or $B$-equivariant modules.
We show the first assertion. By the exactness of ${\mathbb{V}}_{r}$, it only needs to show that ${\mathbb{V}}_{r}{\mathcal{M}}_{x}$ is contained in ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}$, for $x\leq^{\frac{\infty}{2}}t(\nu)$. By forgetting $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-action to $S$-action and forgetting the $B$-equivariant structure, it is equivalent to show that $\pi_{*}{\mathcal{M}}_{x}$ is contained in ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{T,\leq\nu}$. Indeed, by [35, Lem 3.8(2)], for any $w\in W$ and $\lambda\in\Lambda$, in ${\mathcal{O}}_{S}$ we have ${\mathsf{T}}^{-\rho}_{0}M(w\bullet 0+l\lambda)_{S}=M(-\rho+l\lambda)_{S}$. By (4.8) we deduce that $\pi_{*}{\mathcal{M}}_{t(\lambda)w}={\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ (as $T$-equivariant ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-modules), which proves the claim.
We then obtain a natural transformation $\tau_{c}^{\leq\nu}\circ{\mathbb{V}}_{c}\rightarrow{\mathbb{V}}_{c}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$. To show that it is an isomorphism, by forgetting $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-action to $S$-action, it is enough to show the natural isomorphism
(5.7) $\tau_{c}^{\leq\nu}\circ\pi_{*}\rightarrow\pi_{*}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$
By [35, Lem 3.8(2)] again, the module ${\mathsf{T}}_{-\rho}^{0}M(-\rho+l\lambda)_{S}$ is composed by $M(w\bullet 0+l\lambda)_{S}$ with $w\in W$. By (4.9) and Lemma 5.14, the object $\pi^{!}{\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ is composed by ${\mathcal{M}}_{t(\lambda)w}$ with $w\in W$, as $T$-equivariant ${\mathsf{A}}^{b}_{S}$-module. Any object in ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}$ is a quotient of an object composed by ${\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ with $\lambda\leq\nu$. By the exactness of $\pi^{!}$ (it is exact as a functor for $T$-equivariant modules, by (4.9) again), we deduce that $\pi^{!}$ sends ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}$ to ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$. Hence we have a commutative diagram
$({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$$({\mathbb{C}}[{\mathfrak{b}}_{S}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}.$$\scriptstyle{\pi^{!}}$$\scriptstyle{\pi^{!}}$
Now (5.7) follows by taking the left adjunctions. ∎
@@
Let $^{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of $\mathsf{A}$} at $0\in{\mathfrak{t}}/W$. Since the restriction of $\pi^{*}{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ on $\{e\}\times{\mathfrak{b}}\hookrightarrow\widetilde{{\mathfrak{g}}}$ is ${\mathsf{A}}^{b}_{S}$, we can identify
[…]
by induction from $B$ to $G$. The ${\mathsf{A}}^{b}_{S}$-module ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ corresponds to $^{\mathcal{E}}^{\vee}_{\hat{0}}:={\mathcal{E}}^{\vee}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}$ under (5.8). Let $\mathrm{St}=\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{g}}}\widetilde{{\mathfrak{g}}}$ be the Steinberg variety, and let $\widehat{\mathrm{St}}$ be its completion at $0\in{\mathfrak{t}}/W$. Let $\pi_{2}:\widehat{\mathrm{St}}\rightarrow\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ be the projection to the second factor. There is an equivalence
[…]
sending $^{\mathcal{E}}^{\vee}_{\hat{0}}$ to the structure sheaf ${\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{\Delta}}$ of the diagonal $\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ in $\widehat{\mathrm{St}}$.